OpenStudy (anonymous):

Find the angle between the hands of a clock at 5:15.

OpenStudy (anonymous):

at 5 O clock the angle between clocks hand is 180*5/6 =150 degrees after 15 minutes the minute had would have moved , 45 degrees and the hour hand would have moved, 15*60/360=2.5 degrees so the angle between them at 5:15 is 150 - (45-2.5) = 150-45+2.5 =107.5

OpenStudy (anonymous):

Thank you! :D